OpenStudy (anonymous):

During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident?

OpenStudy (anonymous):

t + c = 32 then number of trucks + the number of cars = 32 6t + 4c = 148 each truck has 6 wheels, each car 4 wheels total of 148 Solve: Multiply the 1st equation by -6 -6(t + c = 32) which equals -6t - 6c = -192 6t + 4c = 148 add the two equations... the t's cancel ------------- -2c = -44 c = 22 plug into the first equation t + c = 32 t + 22 = 32 t = 10 22 cars, 10 trucks total of 32 vehicles 22 cars would have 88 wheels 10 trucks would have 60 wheels total of 148 wheels.